Show that the values are equal.
sin^2 60 degrees + cos^2 60 degrees=sin^2 45 degrees + cos^2 45 degrees

Answer by eperette(173)

sin^2 60 degrees + cos^2 60 degrees=sin^2 45 degrees + cos^2 45 degrees
(sin(60))^2 + (cos(60))^2= (sqr(3)/2)^2 + (1/2)^2 = 3/4 + 1/4 = 4/4=1
(sin(45))^2 + (cos(45))^2= (sqr(2)/2)^2 + (sqr (2)/2)^2= (2/4) + (2/4)=4/4=1